OpenStudy (anonymous):

Which of the following represents the radical in simplest form? the square root of b to the fourteenth A. b^7 B. the square root of b^7 C. b^7 and the square root of b

OpenStudy (lalaly):

\[\sqrt{b^{14}}= b^{\frac{14}{2}} = b^7\]